OpenStudy (anonymous):

Is this equation odd, even, or neither? f(x) = -9 x3 + 8x

OpenStudy (anonymous):

Sorry, that's supposed to be Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x^3 + 8x. Whoops.

OpenStudy (freckles):

plug in -x

OpenStudy (freckles):

if f(-x)=f(x) then f is even if f(-x)=-f(x) then f is odd

OpenStudy (freckles):

\[f(x)=-9x^3+8x \\ f(-x)=-9(-x)^3+8(-x)\] can you rewrite (-x)^3?

OpenStudy (anonymous):

and equation is old when the highest power of the independent variable is old

OpenStudy (freckles):

not necessarily true

OpenStudy (freckles):

what if you had f(x)=-9x^3+1 that would be neither

OpenStudy (anonymous):

although they are also some rules guiding it. thanks @freckles

OpenStudy (freckles):

you mean if you have a polynomial and in that polynomial all the powers are odd then the polynomial function is also odd in the example f(x)=-9x^3+1 the powers are 3 and 0

OpenStudy (freckles):

this function is neither because you have not all the powers are odd

OpenStudy (freckles):

and if all the powers are even in the polynomial then the polynomial function is even

OpenStudy (anonymous):

ok. thanks for the correction

OpenStudy (freckles):

but it said to find out algebraically if the function is odd or even or neither and that is why we are doing it the long way above

OpenStudy (anonymous):

Heh.. I suppose this one was a bit self-explanatory. Thanks.